Circular No. 9031

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 F3 (LINEAR) E. Guido, G. Sostero, and P. Camilleri report their recovery of comet

P/2001 MD_{7} (cf. IAUC 7660) on unfiltered CCD exposures obtained remotely with the RAS Observatory 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A., on Mar. 17 and 18 (with ten stacked 120-s images on Mar. 18.49 UT showed a diffuse coma ∼ 15′′ in diameter) and on additional exposures obtained on Mar. 20 remotely with the 0.35-m f/6.7 reflector of Grove Creek Observatory at Trunkey, N.S.W., Australia.

2009 UT  α_{2000}  δ_{2000}  Mag.
Mar. 17.49833  19^{h}43^{m}45.^{s}97  −14^{o}58′54′′.1  18.2
17.50484  19 43 46.60  −14 58 49.5  18.1
18.47991  19 45 35.79  −14 55 18.7  18.9
18.49643  19 45 37.64  −14 55 18.3  18.3
20.74117  19 49 50.08  −14 46 52.0  18.2
20.75159  19 49 51.24  −14 46 49.6  18.1
20.77107  19 49 53.43  −14 46 45.1  18.0

The indicated correction to the predictions on MPC 56804 and the 2008-2009 Comet Handbook is ΔT = +0.01 day. Linked orbital elements by B. G. Marsden and an ephemeris appear on MPEC 2009-F51.

COMETS C/2008 X7–X12 (SOHO) Further to IAUC 9030, additional Kreutz sungrazers have been found

on SOHO website images. C/2008 X7 and C/2008 X8 were very diffuse (peak magnitudes ∼ 8 and 7.5–8, respectively). C/2008 X9 was small and condensed (mag ∼ 7–7.5). C/2008 X10 was tiny and stellar in appearance (mag ∼ 8). C/2008 X11 was small and stellar in appearance (mag ∼ 7) in C3 images, but slightly diffuse with a short, faint, diffuse tail in C2 images. C/2008 X12 was small and stellar in appearance (mag ∼ 6.5) in C3 images, but condensed with a very thin, faint tail in C2 images.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 X7 Dec. 1.329 16^{h}31.^{m}0 −23^{o}40′ C2  MU 2009-E60
C/2008 X8  2.686 16 38.1 −23 55 C2  MA 2009-E60
C/2008 X9  3.604 16 41.2 −24 02 C2  EB 2009-E60
C/2008 X10  3.938 16 42.5 −24 05 C2  RK 2009-E61
C/2008 X11  5.154 16 48.0 −25 51 C3/2 SY 2009-E61
C/2008 X12  8.846 17 02.9 −26 54 C3/2 MU 2009-E61
2009 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green